# Multi-Preference Lambda-weighted Listwise DPO for Dynamic Preference Alignment

### v4

@@ Introduction
Step 1: Polynomial Feature Mapping
where the total number of features is given by:
$N=\binom{d+p}{p}.$ (18)
Step 2: Polynomial Regression
The resulting predictive function
$f(\lambda)=\mathbf{w}^{\top}\phi_{p}(\lambda)$ (20)
Step 3: Grid Construction over the Simplex
Step 4: Softmax-Based Scheduling Distribution
where the temperature parameter $\tau>0$ controls the exploration–exploitation trade-off. Higher values of $f(\lambda^{(j)})$ correspond to better expected performance, allowing the scheduler to prioritize promising $\lambda$ vectors while retaining stochasticity. This yields a flexible, differentiable mechanism for guiding multi-objective alignment through learned preference scheduling.
Step 5: Empirical Instantiation with Five Training Points
$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
We set the polynomial degree to $p=2$, yielding a total of $N=\binom{4+2}{2}=15$ monomial features. Using gradient descent to minimize the squared error over the training set, we obtain the fitted predictive model:
$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
Step 6: Lambda Sampling from Learned Scheduler
We discretize the simplex $\Delta^{4}$ by uniformly sampling $k=10$ candidate $\lambda$ vectors from a Dirichlet$(\boldsymbol{1})$ prior to ensure diverse coverage of the preference space. For each sample $\lambda^{(j)}$, we evaluate its predicted alignment score $s_{j}=f(\lambda^{(j)})$ using the learned polynomial model, and convert these scores into a sampling distribution using softmax:
$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{10}\exp(\tau\cdot f(\lambda^{(l)}))},\quad\text{with }\tau=100.$
This yields a discrete distribution over candidate $\lambda$ vectors that prioritizes those with higher predicted alignment quality. The table below summarizes the 10 sampled vectors, their predicted scores, and their corresponding softmax probabilities:
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $f(\lambda)$ $p(\lambda)$
0.212 0.334 0.245 0.209 0.452 0.095
0.126 0.237 0.131 0.507 0.450 0.075
0.542 0.079 0.256 0.123 0.455 0.117
0.233 0.721 0.020 0.025 0.459 0.185
0.004 0.334 0.281 0.381 0.449 0.067
0.320 0.069 0.513 0.099 0.451 0.082
0.236 0.155 0.358 0.251 0.452 0.093
0.141 0.110 0.701 0.048 0.454 0.110
0.070 0.198 0.388 0.344 0.451 0.081
0.090 0.139 0.306 0.465 0.450 0.095
Table 2: Sampled $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ), their predicted scores $f(\lambda)$, and softmax-based sampling probabilities $p(\lambda)$ with temperature $\tau=100$.
Step 7: Scheduler vs. Uniform Comparison
To assess the overall utility of the learned scheduling distribution, we compare it against a baseline that uses a fixed uniform preference vector $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ for all training examples. Both approaches are applied under identical listwise $\lambda$-DPO training conditions using the LLaMA 3.2B Instruct model.
Figure 4 shows the empirical accuracy across six alignment benchmarks. The learned scheduler achieves a slightly higher average accuracy (0.462 vs. 0.461), validating its effectiveness as a data-driven alternative to static preference weighting.
Beyond average performance, the learned scheduler offers several additional advantages: (i) it introduces sample-level variability by stochastically selecting from a high-quality preference distribution, which can act as an implicit regularizer during training; and (ii) it enables better coverage of the preference simplex, reducing the risk of overfitting to limited preference configurations and helping mitigate modeling bias in the polynomial predictor.
@@ Conclusion
We introduced Multi-Preference Lambda-weighted Listwise DPO, which extends DPO to a listwise setting with controllable $\lambda$ vectors for multi-dimensional alignment. This approach enables flexible, interpretable optimization over preference trade-offs without retraining.
Experiments on 1B–2B scale models show consistent gains over standard DPO, with average accuracy improvements of +1.7% on LLaMA 3.2-1B Instruct and +2.8% on Qwen-1.8 Chat. We further propose a learned scheduler that adaptively samples high-performing $\lambda$ configurations, outperforming uniform baselines while reducing modeling bias. These results highlight the effectiveness of listwise, preference-aware alignment in instruction tuning—especially for efficiently improving smaller models where preference supervision has greater impact, capacity is limited, and retraining is impractical. This is particularly relevant for domain-specific assistants on edge devices, multilingual student models distilled from larger systems, or rapid adaptation via methods like LoRA. In such settings, dynamically adjusting alignment objectives—e.g., helpfulness vs. honesty—without retraining offers both efficiency and controllability.
@@ A.3 Gradient of Lambda-DPO Loss
$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$ (35)
@@
Appendix B B Implementation of _compute_lambda_dpo_loss
This function executes the following procedure:
B.1 Input Tensors and Chunked Computation
  1. 1.
    Load the batch tensors:
    input_ids = inputs["input_ids"]
    attention_mask = inputs["attention_mask"]
    labels = inputs["labels"]
    pi_target = inputs["pi_target"]
  2. 2.
    Assume the batch contains 4 preference dimensions and 4 candidate responses (16 sequences in total). Reshape the tensors so that each preference dimension shares the same 4 unique responses:
    num_pref, num_resp = 4, 4
    B = input_ids.size(0) // (num_pref * num_resp)
    seq_len = input_ids.size(1)
    ids_view = input_ids.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
    mask_view = attention_mask.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
    labels_view = labels.view(B, num_pref * num_resp, seq_len)[:, :num_resp, :].reshape(B * num_resp, seq_len)
  3. 3.
    Choose a chunk size (by default equal to ids_view.size(0)) to compute log-probabilities in smaller batches. Optionally compute reference model probabilities:
    chunk_size = self.lambda_dpo_chunk_size or ids_view.size(0)
    seq_log_probs_list = []
    ref_seq_log_probs_list = []
    for start in range(0, ids_view.size(0), chunk_size):
    end = start + chunk_size
    chunk_ids = ids_view[start:end]
    chunk_mask = mask_view[start:end]
    chunk_labels = labels_view[start:end]
    logits = model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
    seq_lp, valid_len = get_batch_logps(logits, chunk_labels, self.label_pad_token_id)
    seq_log_probs_list.append(seq_lp / valid_len)
    with torch.no_grad():
    if self.finetuning_args.use_ref_model and self.ref_model is not None:
    ref_logits = self.ref_model(input_ids=chunk_ids, attention_mask=chunk_mask).logits
    ref_lp, ref_len = get_batch_logps(ref_logits, chunk_labels, self.label_pad_token_id)
    ref_lp = ref_lp / ref_len
    else:
    ref_lp = torch.zeros_like(seq_lp)
    ref_seq_log_probs_list.append(ref_lp)
  4. 4.
    Concatenate the chunk results and share each set of four log-probabilities across the preference dimensions:
    seq_log_probs = torch.cat(seq_log_probs_list, dim=0)
    ref_seq_log_probs = torch.cat(ref_seq_log_probs_list, dim=0)
    seq_log_probs = seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
    ref_seq_log_probs = ref_seq_log_probs.view(B, num_resp).unsqueeze(1).expand(B, num_pref, num_resp).reshape(-1)
B.2 Reference Model and Relative Reward
  1. 5.
    Compute the relative reward and reshape back to $(B,4,4)$:
    r_theta = self.beta * (seq_log_probs - ref_seq_log_probs)
    r_theta = r_theta.view(B, num_pref, num_resp)
    pi_target = pi_target.view(B, num_pref, num_resp)
B.3 Listwise Loss by Preference Dimension
  1. 6.
    For each preference axis $i\in\{1,2,3,4\}$, compute the listwise loss:
    listwise_losses = []
    for i in range(num_pref):
    r = r_theta[:, i, :]
    p = pi_target[:, i, :] + 1e-10
    log_softmaxed = r - torch.logsumexp(r, dim=1, keepdim=True)
    loss = -(p * log_softmaxed).sum(dim=1).mean()
    listwise_losses.append(loss)
B.4 Lambda Weights and Final Loss
  1. 7.
    The current implementation equally weights the preference dimensions (no random sampling of $\lambda$):
    num_losses = len(listwise_losses)
    weights = torch.ones(num_losses, device=input_ids.device)
    weights /= weights.sum()
    final_loss = sum(w * l for w, l in zip(weights, listwise_losses))
    if return_outputs:
    return final_loss, {"loss_components": listwise_losses}
    return final_loss
Appendix C: Humorous Overview of the Added Classes
C.1 UltrafeedbackDatasetConverter
Input
A single UltraFeedback record. It contains:
  • a user prompt (instruction)
  • multiple completions, each with four ratings: helpfulness, honesty, instruction_following, and truthfulness.
Output
A dictionary with the prompt and, for every dimension, the raw scores plus a preference vector.
Example
Input:
instruction = "Explain photosynthesis."
completions = [
{"response": "...", "annotations": {"helpfulness": {"Rating": 4}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 2}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 5}}},
{"response": "...", "annotations": {"helpfulness": {"Rating": 3}}}
Output (simplified):
"_prompt": [{"role": "user", "content": "Explain photosynthesis."}],
"helpfulness": [
{"model": "?", "content": "...", "score": 4},
"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22]}
C.2 ListwiseDatasetProcessor
Input
A dictionary produced by the converter that contains _prompt, the lists of responses for each dimension, and _pi_target.
Output
A flattened structure ready for training: for every sample there are 16 sequences of token IDs (one per response per dimension), matching labels and attention masks, and the list of probabilities.
Example
Input:
"_prompt": ...,
"helpfulness": [
{"content": "A"}, {"content": "B"}, {"content": "C"}, {"content": "D"}
"_pi_target": {"helpfulness": [0.24, 0.09, 0.45, 0.22], ...}
Output (illustrative shapes):
"input_ids": [seq1, seq2, ..., seq16], # each seq corresponds to a prompt+response
"labels": [...], # same layout as input_ids
"attention_mask": [...], # masks per sequence
"pi_target": [0.24, 0.09, 0.45, 0.22, ...] # flattened list of probabilities
C.3 ListwiseDataCollatorWithPadding
Input
A list of processed samples where each sample holds 16 sequences of token IDs and the flattened probability list.
Output
A padded batch containing tensors input_ids, attention_mask, labels, and an accompanying tensor of preference values aligned with the responses.
Example
Input:
# List of samples where each sample holds 16 sequences and
# a corresponding 16-element probability list
[example1, example2, ..., example_B] # B = batch size
Output:
"input_ids": Tensor of shape [B * 16, L],
"attention_mask": Tensor of shape [B * 16, L],
"labels": Tensor of shape [B * 16, L],
"pi_target": Tensor of shape [B * 16]
These three classes transform UltraFeedback data into tensors suitable for Lambda-weighted DPO training, while preserving the four-response structure and converting ratings into softmax preference vectors.
This function performs sequence probability calculation, reference model comparison, listwise cross-entropy over preference dimensions, and $\lambda$-weighted aggregation. It fully realizes the lambda-DPO loss objective described in the paper’s methodology section.
Appendix C D Implementation of _compute_ultra_dpo_loss
This routine computes the DPO loss when training on the UltraFeedback dataset, where each example contains four candidate responses with human preference scores. It proceeds as follows:
D.1 Input Aggregation
  1. 1.
    Load the tensors input_ids, attention_mask, labels, and pi_target.
  2. 2.
[…]
  3. 3.
    Apply uniform weights across the four preference dimensions to obtain an aggregated distribution for each example:
    $\text{aggregated}=\sum_{i=1}^{4}\tfrac{1}{4}\,\pi^{*(i)}.$
D.2 Selecting Best and Worst Responses
  1. 4.
    Determine the best response index via $\arg\max$ over aggregated and the worst index via $\arg\min$.
  2. 5.
    Gather the corresponding best and worst sequences from the original batch for further processing.
D.3 Log-Probability Computation
  1. 6.
    Pass the selected sequences through the model, compute token log-probabilities, and average over tokens to obtain sequence log-probabilities.
  2. 7.
    If a reference model is provided, compute its log-probabilities in the same way; otherwise use zero tensors as the reference scores.
D.4 Final DPO Loss
  1. 8.
    Split the policy and reference log-probabilities into best and worst halves.
  2. 9.
    Invoke compute_preference_loss on these values to obtain per-example DPO losses, then average across the batch:
    $\mathcal{L}_{\mathrm{DPO}}=\frac{1}{B}\sum_{b=1}^{B}\mathrm{loss}(y^{\mathrm{best}}_{b},\,y^{\mathrm{worst}}_{b}).$
Appendix D E Accuracy Comparisons on UltraFeedback Benchmarks with 1B model

### v5

@@ Introduction
Performance-Driven Scheduler Design
We propose a performance-driven scheduler to improve multi-objective alignment by dynamically selecting preference vectors $\lambda$ based on their estimated effectiveness. Instead of fixing a single $\lambda$, the scheduler learns a distribution over $\lambda$ configurations that reflect observed performance.
The scheduler fits a polynomial regression model to past $(\lambda,\text{accuracy})$ pairs and uses it to score new candidates sampled from a Dirichlet prior over the simplex. A softmax over predicted scores then defines sampling probabilities, balancing exploitation of high-performing regions and exploration. This adaptive strategy promotes training diversity and improves robustness.
Empirical Comparison with Uniform Baseline
As shown in Figure 12, the learned scheduler achieves slightly better average accuracy (0.462 vs. 0.461) across six evaluation benchmarks when compared with a static baseline that uses the uniform configuration $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$. While the performance gain is modest, the scheduler enhances robustness and reduces overfitting by continuously adapting to observed training signals.
Importantly, this framework supports flexible preference adaptation and can be extended to personalized or task-specific settings without requiring retraining.
@@
Extension to LLaMA-8B
To assess the scalability of $\lambda$-DPO, we extend our experiments to the LLaMA-8B model. Across six standard benchmarks, $\lambda$-DPO yields an average relative improvement of approximately 0.3% over DPO, compared to a much larger average gain of 1.7% on the LLaMA-1B model.
Figure 6: Accuracy comparison between DPO and $\lambda$-DPO on LLaMA-8B across six benchmarks.
As shown in Figure 6, $\lambda$-DPO achieves better alignment on the 1B model, with notably larger gains on challenging tasks such as GSM8K. This suggests that smaller models benefit more from listwise multi-preference supervision, likely because alignment signals have a stronger relative effect when model capacity is limited.
Moreover, training the 1B model only requires around 20GB of GPU memory, making it both efficient and accessible for academic labs, startups, or edge-focused applications. This low resource footprint enables rapid experimentation and deployment on commodity hardware, without sacrificing alignment quality. As such, $\lambda$-DPO is well-suited for scalable, real-world use cases where cost, speed, and controllability are critical.
@@ Conclusion
In this paper, we propose Multi-Preference Lambda-weighted Listwise DPO ($\lambda$-DPO) to enable efficient, controllable, and fine-grained alignment of language models with diverse human preferences.
For robust and flexible alignment, we generalize pairwise supervision to listwise preference modeling, allowing the model to learn from ranked completions and capture nuanced preference structures. To support personalized and multi-objective alignment, we introduce a simplex-weighted $\lambda$ vector that dynamically balances objectives such as helpfulness, honesty, and fluency—without retraining.
To enhance generalization and robustness, we incorporate a learned scheduler that adaptively samples high-performing $\lambda$ configurations during training, improving diversity while mitigating modeling bias. Additionally, the method is designed to be resource-efficient, requiring only 20GB of GPU memory, making it practical for deployment in constrained scenarios such as on-device assistants and educational tools.
Experimental results on LLaMA 3.2-1B and Qwen-1.8B demonstrate that $\lambda$-DPO achieves consistent improvements over standard DPO, with average accuracy gains of +1.7% and +2.8%, respectively, across standard alignment benchmarks.
@@ A.3 Gradient of Lambda-DPO Loss
$=-\mathbb{E}_{x,\lambda}\left[\sum_{i}\left(p^{\lambda}(y_{i}\mid x)-P_{\theta}(y_{i}\mid x)\right)\nabla_{\theta}\log\pi_{\theta}(y_{i}\mid x)\right]$
@@
Appendix B B Experiment Setup
B.1 Experimental Objectives and Comparison Targets
We consider two levels of comparison:
Model comparison.
We experiment with two strong open-source language models—LLaMA and Qwen. These models differ in architecture, pretraining corpora, and tokenization schemes, providing a robust testbed for evaluating generalizability.
Method comparison.
We compare DPO and $\lambda$-DPO. For each method-model pair, we perform independent hyperparameter tuning (e.g., learning rate, batch size, $\beta$) to ensure comparisons are fair and reflect the best achievable performance under comparable resource constraints.
B.2 Training Settings
Inspired by the SimPO baseline, we adopt a unified training pipeline:
  • Finetuned-Base: The model is first SFT-trained on UltraChat-200k, then optimized on UltraFeedback via DPO or $\lambda$-DPO to reflect realistic alignment scenarios.
B.3 Evaluation Benchmarks and Metrics
We evaluate model performance using the lm-eval-harness (Gao et al. 2024) framework, which reports accuracy on diverse benchmarks covering reasoning, factual correctness, and commonsense understanding.
  • MMLU: 57 subjects, 15K questions assessing academic knowledge (Hendrycks et al. 2020).
  • ARC: 1.1K grade-school science questions requiring reasoning (Clark et al. 2018).
  • HellaSwag: 10K commonsense sentence completions with distractors (Zellers et al. 2019).
  • TruthfulQA: 817 questions testing factuality and hallucination (Lin, Hilton, and Evans 2021).
  • Winograd: 273 examples probing coreference and syntax (Levesque, Davis, and Morgenstern 2012).
  • GSM8K: 8.5K arithmetic reasoning problems (Cobbe et al. 2021).
B.4 Ablation: Impact of Lambda Weighting
  • Uniform $\lambda$: Equal weights across all dimensions ($\lambda=(1/m,\dots,1/m)$), simulating balanced multi-objective aggregation.
  • Unimodal $\lambda$: One-hot vectors (e.g., $(1,0,\dots)$) isolating single-dimension optimization without interpolation.
This comparison helps isolate the contribution of $\lambda$-based preference interpolation within the listwise setting.
Appendix C C Accuracy Comparisons on UltraFeedback Benchmarks with 1B model
@@ C.5 Unimodal Ablation
Appendix D D Performance-Driven Lambda Scheduling via Polynomial Modeling
D.1 Polynomial Feature Mapping
where the total number of features is given by:
$N=\binom{d+p}{p}.$ (32)
D.2 Polynomial Regression
The resulting predictive function
$f(\lambda)=\mathbf{w}^{\top}\phi_{p}(\lambda)$ (34)
D.3 Grid Construction over the Simplex
D.4 Softmax-Based Scheduling Distribution
where the temperature parameter $\tau>0$ controls the exploration–exploitation trade-off.
D.5 Empirical Instantiation with Five Training Points
$\lambda^{(1)}$ $=[1,0,0,0],\quad y^{(1)}=0.4593$
$\lambda^{(2)}$ $=[0,1,0,0],\quad y^{(2)}=0.4598$
$\lambda^{(3)}$ $=[0,0,1,0],\quad y^{(3)}=0.4543$
$\lambda^{(4)}$ $=[0,0,0,1],\quad y^{(4)}=0.4548$
$\lambda^{(5)}$ $=[0.25,0.25,0.25,0.25],\quad y^{(5)}=0.4612$
We set $p=2$, yielding $N=\binom{4+2}{2}=15$ features. Using gradient descent to minimize squared error, we obtain:
$+0.01\lambda_{3}^{2}+0.00\lambda_{3}\lambda_{4}+0.01\lambda_{4}^{2}$
D.6 Lambda Sampling from Learned Scheduler
We sample $k=10$ $\lambda$ vectors from Dirichlet(1,1,1,1)$1$(1,1,1,1)( 1 , 1 , 1 , 1 ) and compute predicted scores $f(\lambda^{(j)})$, converting them via softmax ($\tau=100$):
$p(\lambda^{(j)})=\frac{\exp(\tau\cdot f(\lambda^{(j)}))}{\sum_{l=1}^{10}\exp(\tau\cdot f(\lambda^{(l)}))}$
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $f(\lambda)$ $p(\lambda)$
0.212 0.334 0.245 0.209 0.452 0.095
0.126 0.237 0.131 0.507 0.450 0.075
0.542 0.079 0.256 0.123 0.455 0.117
0.233 0.721 0.020 0.025 0.459 0.185
0.004 0.334 0.281 0.381 0.449 0.067
0.320 0.069 0.513 0.099 0.451 0.082
0.236 0.155 0.358 0.251 0.452 0.093
0.141 0.110 0.701 0.048 0.454 0.110
0.070 0.198 0.388 0.344 0.451 0.081
0.090 0.139 0.306 0.465 0.450 0.095
Table 2: Sampled $\lambda$ vectors, their predicted scores $f(\lambda)$, and softmax sampling probabilities $p(\lambda)$ at temperature $\tau=100$.
D.7 Scheduler vs. Uniform Comparison
We compare our learned scheduler with a uniform $\lambda=[\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4}]$ baseline using LLaMA 3.2B Instruct. The learned scheduler achieves slightly better average accuracy (0.462 vs. 0.461), and offers regularization via stochastic $\lambda$ sampling, enhancing robustness.
Figure 12: Comparison of learned scheduler vs. fixed uniform $\lambda$ across six alignment benchmarks. The learned scheduler yields improved accuracy and diversity.